OpenStudy (anonymous):

Calculus question: Use Part 1 of FTC to find the derivative of the function

OpenStudy (anonymous):

\[g(x)=\int\limits_{4}^{x}e ^{4t ^{2}-4} dt\] gprime(x)=

OpenStudy (anonymous):

Just replace the x with the variable t

OpenStudy (anonymous):

It's that easy

OpenStudy (anonymous):

wow...i feel dumb >_<

OpenStudy (anonymous):

\[\huge g'(x) = e^{4x^2-4}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Np :)

OpenStudy (anonymous):

The fundamental theorem of calc part 1 states this, If f is continuous on [a,b] then \[g(x) = \int\limits_{a}^{x} f(t) dt~~~~a \le x \le b\] is continuous on [a,b] differentiable on (a,b) and \[g'(x) = f(x)\]

OpenStudy (anonymous):

Basically saying plug in x for t haha.

OpenStudy (anonymous):

Or what ever the variable is :P